# Flatten Anything: Unsupervised Neural Surface Parameterization

### v1

@@ Abstract
Surface parameterization plays an essential role in numerous computer graphics and geometry processing applications. Traditional parameterization approaches are designed for high-quality meshes laboriously created by specialized 3D modelers, thus unable to meet the processing demand for the current explosion of ordinary 3D data. Moreover, their working mechanisms are typically restricted to certain simple topologies, thus relying on cumbersome manual efforts (e.g., surface cutting, part segmentation) for pre-processing. In this paper, we introduce the Flatten Anything Model (FAM), an unsupervised neural architecture to achieve global free-boundary surface parameterization via learning point-wise mappings between 3D points on the target geometric surface and adaptively-deformed UV coordinates within the 2D parameter domain. To mimic the actual physical procedures, we ingeniously construct geometrically-interpretable sub-networks with specific functionalities of surface cutting, UV deforming, unwrapping, and wrapping, which are assembled into a bi-directional cycle mapping framework. Compared with previous methods, our FAM directly operates on discrete surface points without utilizing connectivity information, thus significantly reducing the strict requirements for mesh quality and even applicable to unstructured point cloud data. More importantly, our FAM is fully-automated without the need for pre-cutting and can deal with highly-complex topologies, since its learning process adaptively finds reasonable cutting seams and UV boundaries. Extensive experiments demonstrate the universality, superiority, and inspiring potential of our proposed neural surface parameterization paradigm. The code will be publicly available.
@@ 2.1 Traditional Parameterization Approaches
Within the computer graphics and geometry processing communities, mesh parameterization [10, 34] is a fundamental and long-standing task, which has been extensively investigated owing to its diverse downstream applications, such as remeshing, morphing, editing, texture mapping, and compression. Early approaches typically formulate this task as a Laplacian problem, where boundary points are anchored to a pre-determined convex 2D curve [7, 9], tackled by sparse linear system solvers. Such a linear strategy is appreciated for its simplicity, efficiency, and the guarantee of bijectivity. However, the rigidity of fixing boundary points in the parameter domain usually leads to significant distortions. In response to these issues, free-boundary parameterization algorithms [33, 19] have been proposed, offering a more flexible setting by relaxing boundary constraints. Although these approaches bring certain improvements, they often struggle to maintain global bijectivity. More recent state-of-the-art approaches [29, 36] shift towards minimizing simpler proxy energies by alternating between local and global optimization phases, which not only accelerate convergence but also enhance the overall parameterization quality. Despite continuous advancements, the working mechanisms of all these approaches fundamentally rely on the connectivity information inherent in mesh structures, casting doubt on their applicability to unstructured point clouds. Moreover, mesh-oriented parameterization approaches typically assume that the inputs are well-behaved meshes possessing relatively regular triangulations. When faced with input meshes of inferior quality characterized by irregular triangulations and anomalies, remeshing becomes an indispensable procedure. Additionally, how to deal with non-ideal geometries and complex topologies (e.g., non-manifolds, multiply-connected components, thin structures) has always been highly challenging.
@@ 3.1 Bi-directional Cycle Mapping
$[\mathbf{\hat{P}};\mathbf{\hat{P}_{n}}]=\mathcal{M}_{w}(\mathbf{\hat{Q}})=\xi_{w}^{\prime\prime}([\xi_{w}^{\prime}(\mathbf{\hat{Q}});\mathbf{\hat{Q}}]),$ (2)
@@ 4 Experiments
We collected a series of 3D surface models with different types and complexities of geometric and/or topological structures for experimental evaluation. We made qualitative and quantitative comparisons with SLIM [29], a state-of-the-art and widely-used mesh parameterization algorithm. Additionally, we also compared our FAM with FBCP-PC [4] for unstructured and unoriented (i.e., without normals) 3D point cloud parameterization. Finally, we conducted ablation studies to demonstrate the effectiveness of the proposed bi-directional cycle mapping framework. Due to page limitations, detailed technical implementations and additional experimental results are presented in our Appendix.
@@ Appendix A Appendix
In this appendix, we presented the technical implementation details of our approach. Additionally, we provided other aspects of comparisons with the state-of-the-art mesh parameterization algorithm SLIM [29] to facilitate more comprehensive understandings of our advantages.
@@
In our main manuscript, our experimental comparisons with SLIM focus on 3D models with open boundaries and/or disk topologies, which SLIM can be applied on. Here, we made additional efforts to conduct comparisons with SLIM on the non-disk-type 3D surface model of nefertiti by manually specifying an optimal cutting seam for SLIM to run. Note that this is a quite unfair experimental setting, because finding optimal cuts is known to be a critical yet complicated task. As illustrated in Figure 7, our approach still achieves highly competitive performances compared with SLIM even under such an unfair comparison setup. In addition to parameterization quality, we further compared the time costs of our FAM and SLIM for training and optimization. Here, the official code of SLIM runs on the Intel(R) Core(TM) i7-9700 CPU. Table 4 lists the time costs for different testing models. Note that, since our FAM operates on surface-sampled points and we uniformly used the same number of input points (i.e., $N$), the time costs of our FAM basically maintain unchanged for different testing models. It turns out that our approach achieves satisfactory learning efficiency.

### v2

@@ Abstract
Surface parameterization plays an essential role in numerous computer graphics and geometry processing applications. Traditional parameterization approaches are designed for high-quality meshes laboriously created by specialized 3D modelers, thus unable to meet the processing demand for the current explosion of ordinary 3D data. Moreover, their working mechanisms are typically restricted to certain simple topologies, thus relying on cumbersome manual efforts (e.g., surface cutting, part segmentation) for pre-processing. In this paper, we introduce the Flatten Anything Model (FAM), an unsupervised neural architecture to achieve global free-boundary surface parameterization via learning point-wise mappings between 3D points on the target geometric surface and adaptively-deformed UV coordinates within the 2D parameter domain. To mimic the actual physical procedures, we ingeniously construct geometrically-interpretable sub-networks with specific functionalities of surface cutting, UV deforming, unwrapping, and wrapping, which are assembled into a bi-directional cycle mapping framework. Compared with previous methods, our FAM directly operates on discrete surface points without utilizing connectivity information, thus significantly reducing the strict requirements for mesh quality and even applicable to unstructured point cloud data. More importantly, our FAM is fully-automated without the need for pre-cutting and can deal with highly-complex topologies, since its learning process adaptively finds reasonable cutting seams and UV boundaries. Extensive experiments demonstrate the universality, superiority, and inspiring potential of our proposed neural surface parameterization paradigm. Our code is available at https://github.com/keeganhk/FlattenAnything.
@@ 2.1 Traditional Parameterization Approaches
Early approaches formulate mesh parameterization as a Laplacian problem, where boundary points are anchored to a pre-determined convex 2D curve [7, 9], tackled by sparse linear system solvers. Such a linear strategy is appreciated for its simplicity, efficiency, and the guarantee of bijectivity. However, the rigidity of fixing boundary points in the parameter domain usually leads to significant distortions. In response to these issues, free-boundary parameterization algorithms [33, 19] have been proposed, offering a more flexible setting by relaxing boundary constraints. Although these approaches bring certain improvements, they often struggle to maintain global bijectivity. More recent state-of-the-art approaches [29, 35] shift towards minimizing simpler proxy energies by alternating between local and global optimization phases, which not only accelerate convergence but also enhance the overall parameterization quality. OptCuts [16] explores joint optimization of surface cutting and mapping distortion. Despite the continuous advancements, the working mechanisms of all these approaches fundamentally rely on the connectivity information inherent in mesh structures, casting doubt on their applicability to unstructured point clouds. Moreover, mesh-oriented parameterization approaches typically assume that the inputs are well-behaved meshes possessing relatively regular triangulations. When faced with input meshes of inferior quality characterized by irregular triangulations and anomalies, remeshing becomes an indispensable procedure. Additionally, how to deal with non-ideal geometries and complex topologies (e.g., non-manifolds, multiply-connected components, thin structures) has always been highly challenging.
@@ 3.1 Bi-directional Cycle Mapping
$[\mathbf{\hat{P}};\mathbf{\hat{P}^{n}}]=\mathcal{M}_{w}(\mathbf{\hat{Q}})=\xi_{w}^{\prime\prime}([\xi_{w}^{\prime}(\mathbf{\hat{Q}});\mathbf{\hat{Q}}]),$ (2)
@@ 4 Experiments
We collected a series of 3D surface models with different types and complexities of geometric and/or topological structures for experimental evaluation. We made qualitative and quantitative comparisons with SLIM [29], a state-of-the-art and widely-used mesh parameterization algorithm. Additionally, we also compared our FAM with FBCP-PC [4] for unstructured and unoriented (i.e., without normals) 3D point cloud parameterization. Finally, we conducted ablation studies to demonstrate the effectiveness of the proposed bi-directional cycle mapping framework and the necessity of Cut-Net, together with different aspects of verifications for comprehensively understanding the characteristics of our approach. Due to page limitations, detailed technical implementations and additional experimental results are presented in our Appendix
@@
In addition, we conducted quantitative evaluations of self-intersection. Given a triangular mesh, we check if any pair of triangles overlaps in the UV space, and then measure the proportion of overlapped pairs to the total amount of triangle pairs. As reported in Table 8, self-intersection are inevitable both in FAM and SLIM, but the proportions of self-intersected triangles are very small. In practice, it is not hard to apply post-processing refinement to slightly adjust the distribution of UV coordinates to further relieve or even eliminate self-intersection issues. For point cloud parameterization, we evaluated our robustness to noises in Figure 8, where we can observe that FAM shows stable performances to noisy input conditions (with 1%, 2%, and 4% Gaussian noises for point position perturbation).
[…]
Figure 7: Ablation study on the necessity of Cut-Net: (a) input meshes; (b) results obtained from our full FAM architecture; (c) results obtained by removing the Cut-Net component.
Testing Models FAM SLIM
Open-Surface Models (as in Figure 3) 0.156% 0.133%
Higher-Genus Models (as in Figure 4) 0.204% N/A
Table 4: Quantitative self-intersection metrics of our parameterization results.
Figure 8: Applying FAM to point clouds added with different levels of Gaussian noises.
(a) Highly challenging stress tests of our FAM.
(b) Failure cases of complicated CAD models.
Figure 9: (a) input meshes; (b) UV maps; (c): texture mappings and learned cutting seams.
Finally, we performed stress tests on a Hilbert-space-filling-shaped cylinder model in Figure 9(a). It is observed that our FAM obtains the basically optimal solution. Still, as shown in Figure 9(b), processing the highly-complicated ShapeNet-style CAD model with rich interior structures and many multi-layer issues shows inferior UV unwrapping quality. Although our learned cutting seams are generally reasonable and texture mapping looks relatively regular from outside, it is hard to deal with the complicated interior structures.
@@ Appendix A Appendix
A.1 Implementation Details
@@
A.2 Discussions about Global and Multi-Chart Surface Parameterization
Over the years, global surface parameterization has continuously been the mainstream direction of research, since it potentially achieves smooth transitions and uniform distribution of mapping distortion across the entire surface, while multi-chart parameterization typically introduces discontinuities along patch boundaries, causing more obvious visual artifacts for texture mapping (perhaps the most important application scenario in the graphics field) and bringing additional difficulties in many other shape analysis tasks (e.g., remeshing, morphing, compression). For multi-chart parameterization, it is worth emphasizing that only obtaining chart-wise UV maps is not the complete workflow. We need to further perform chart packing to compose the multiple UV domains, without which the actual usefulness can be largely weakened. However, packing is known as a highly non-trivial problem, which is basically skipped in recent neural learning approaches, such as DiffSR and Nuvo. Still, local parameterization does have its suitable application scenarios for processing highly-complicated surfaces such as typical ShapeNet-style CAD meshes with rich interior structures and severe multi-layer issues, and per-chart distortion can be reduced since the geometry and topology of cropped surface patches become simpler.
A.3 Comparison with SLIM on Non-Disk-Type Surfaces with Manually-Specified Cuts
In the main manuscript, our experimental comparisons with SLIM focus on 3D models with open boundaries and/or disk topologies, on which SLIM can be applied. Here, we made additional efforts to conduct comparisons with SLIM on the non-disk-type 3D surface model of nefertiti by manually specifying an optimal cutting seam for SLIM to run. Note that this is a quite unfair experimental setting, because finding optimal cuts is known to be a critical yet complicated task. As illustrated in Figure 10, our approach still achieves highly competitive performances compared with SLIM even under such an unfair comparison setup.
A.4 Running Efficiency
In addition to parameterization quality, we further compared the time costs of our FAM and SLIM for training and optimization. Here, the official code of SLIM runs on the Intel(R) Core(TM) i7-9700 CPU. Table 5 lists the time costs for different testing models. Note that, since our FAM operates on surface-sampled points and we uniformly used the same number of input points (i.e., $N$), the time costs of our FAM basically maintain unchanged for different testing models. It turns out that our approach achieves satisfactory learning efficiency.
Figure 11: Results of Nuvo: (a) input meshes (different colors denote chart assignment); (b) chart-wise UV maps.
Figure 12: Results of OptCuts: (a) input meshes; (b) UV maps.
A.5 Comparison with Nuvo [36]
We provided typical parameterization results of Nuvo using a third-party implementation available at https://github.com/ruiqixu37/Nuvo, as presented in Figure 11. We can observe that its chart assignment capability is still not stable. It often occurs that some spatially-disconnected surface patches are assigned to the same chart. Moreover, the critical procedure of chart packing is actually ignored in Nuvo. Directly merging the rectangular UV domains is not a valid packing. A real-sense packing should adaptively adjust the positions, poses, and sizes of each UV domain via combinations of translation, rotation, and scaling.
A.6 Comparison with OptCuts [16]
We provided typical parameterization results of OptCuts, as presented in Figure 12. Comparatively, our neural parameterization paradigm still shows advantages in terms of flexibility (not limited to well-behaved meshes; applicable to point clouds), convenience (exploiting GPU parallelism; much easier for implementing and tuning), and parameterization smoothness (not limited to mesh vertices). Although OptCuts is able to jointly obtain reasonable surface cuts and UV unwrapping results, its performances are generally sub-optimal compared with ours.